# Astrophysical Black holes: An Explanation for the Galaxy Quenching

### v2

@@ I Introduction
In this paragraph, we outline the basic picture set out in the present study. The long-term quenching of galaxies is primarily attributed to a dense, compact object situated at the galactic nucleus[41]. Conventional black holes possess event horizon, limiting matter accretion up to a boundary which is set at the “Innermost stable Circular Orbit” (ISCO) and the consequent upper bound on the generation of ultra-high-energy particles. The cosmic censorship conjecture[42, 43] remains a foundational yet unproven and not even properly formulated, assumption in black hole physics. While it holds in simplified models such as spherically symmetric dust collapse, a wide range of studies on more general collapse scenarios—including those involving inhomogeneous matter fields, pressure, and rotation—have shown that naked singularities can emerge as legitimate solutions to Einstein’s equations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Despite multiple attempts, no rigorous mathematical proof exists for any version of cosmic censorship, and the conjecture’s vague formulation continues to hinder progress. In this light, the theoretical exclusion of naked singularities is no longer viable. Since the formation of such singularities is a generic outcome in many physically reasonable collapse models, the focus now necessarily shifts toward their observational signatures. In this work, we consider the astrophysical implications of naked singularities—viewed as ultra-compact, massive objects without event horizons—and explore their potential impact on galaxy evolution, particularly in comparison to traditional black holes. To address this, we propose here models of “astrophysical black holes” (ABHs) that are devoid of an event horizon (EH), hence characterized by a huge amount of matter confined within an extremely compact region, a situation which permits the disk of accreting matter to extend all the way to the centre and yield extremely high energy densities and temperatures. This access to the ultra-strong gravity regime facilitates generation of extremely energetic particles, thus contributing to the long-term quenching of star formation in the host galaxy and its halo region. Additionally, we introduce models of “stellar-mass astrophysical black holes” (StMABHs) which can arise from gravitational collapse of massive stars once they run out of their internal nuclear fuel. These StMABHs, prior to getting enveloped by the event horizon, create high-density regions capable of emitting very high-energy particles over short time scales. Such a situation may lead to short-term episodes of (localised) intra-galactic quenching. Potential observational signatures associated with the StMABHs are also pointed out here. Section II provides a brief overview of the recent observations related to this phenomenon, made with the James Webb Space Telescope (JWST). Section III touches upon the role SMBHs play in the galaxy quenching process. Section IV presents some salient features of our proposed models involving astrophysical black holes (ABH), in the context of long- and short-term quenching. Our conclusions are summarised in Section V.
@@ IV.1 A plausible scenario for long-term quenching
[…]
@@
[…]
@@ IV.3 The Role of Radiative and Mechanical Efficiency
Black holes with super-Eddington accretion would operate in a radiatively efficient, ‘quasar’ mode. In contrast, BHs with highly sub-Eddington accretion operate in a radiatively inefficient, ‘radio’ mode. Despite the lower radiative efficiency, this ‘radio’ mode generates substantial mechanical luminosity, driving winds and/or collimated jets[94]. The magnitude of the Eddington ratio determines the transition between the two accretion (i.e., feedback) modes. We recall the simplified treatment by Silk and Rees[74], where they estimated the effect of the accretion disk winds on the galaxy. The galaxy was assumed to be an isothermal sphere having a gas fraction, $f_{gas}$ and a density $\rho=v^{2}/2\pi Gr^{2}$, where $v$ is the velocity dispersion. The wind can propel the interstellar gas out of the galaxy with a constant velocity,
@@
Here, the wind luminosity is the fraction $f_{w}$ of the Eddington luminosity. The fraction $f_{w}$ is dependent on the radiation efficiency $\epsilon$ as,
@@
where $\gamma\equiv f_{gas}/32\pi^{3}f_{w}$. In theory then, a black hole could expel all the material out of its host galaxy, once its mass surpasses $\sim 10^{7}M_{\odot}$(for $\gamma\sim 1$ $f_{w}\sim 0.01$)(however, these arguments hold for early galaxies).
@@
Here, the superscipt $\dot{}$ represents the partial derivative with respect to $t$ and $g$ represents the determinant of the metric, $g=-g_{tt}g_{rr}g_{\theta\theta}g_{\phi\phi}$. $r_{in}$ is the innermost stable circular orbit, which can be found by requiring,
@@
A more practical measure is the luminosity, i.e. the energy that reaching an observer at infinity $\mathcal{L}_{\infty}$, per unit time. By integrating $d\mathcal{L}_{\infty}/dlnr=4\pi r\sqrt{-g}E\mathcal{F}$ with $lnr$, we can calculate the net luminosity $\mathcal{L}_{\infty}$ observed at infinity. As the disc is truncated at the ISCO in the BH case, the efficiency is only $\sim 6\%$ as $\epsilon=1-E_{ISCO}=\mathcal{L}_{\infty}/\dot{m}c^{2}$. In the case of ABHs (see, JMN1 & JMN2), the accretion disk can extend up to the central high-density entity. However, the singularity is not yet formed, so the accretion disk extends up to the boundary of the central ultradense object, and the energy of the innermost circular orbit will then always be lesser than the $E_{ISCO}$. This essentially means that, as the collapse proceeds, the radiative efficiency of the ABH rises. When the singularity has formed in the asymptotic time, the stable circular orbits can extend right up to the singularity. The radiative efficiency will then approach $100\%$. The $\gamma$ of the Eqn. (5) will be $\sim 17$ in that case, and we require $17$ times more mass than the typical BH, in order to expel all the material from the host galaxy. This estimate is meant for the feedback due to mechanically generated winds from the accretion disk. The calculation assumes that the mechanical efficiency is reversibly proportional to the radiative efficiency. The radiative efficiency of ABH being very high, it would have a lower efficiency for mechanical energy injection and hence required to be more massive. However, when the winds are accelerated by the disc radiation, the ABH will be able to generate high-energy winds with greater efficiency. The next subsection discusses this in detail with the example of the hydrodynamical simulation for a naked singularity spacetime. In the BH case, the radiation-dominated (quasar) mode sets in only when the accretion rate becomes very high. In contrast, an ABH can produce very efficient radiation-driven winds even when the accretion rate is very low.
@@ IV.4 Wind Generation
with ionization rate $\chi_{e}=1$, Eddington luminosity $L_{E}=4\pi GcMm_{p}/\sigma_{T}$ (here $m_{p}$ is the proton mass), and the disk luminosity,
@@
In the units of $2GM/c^{2}$, the ISCO for the Schwarzschild black hole is located at $r_{0}=3$. However, the ISCO for the ABH without EH models discussed in the previous section would be very close to the central ultra-dense object. If we assume the ultra compact object to be of the size of the order of Planck length, then ISCO for a few solar mass ABH without EH will be at $r_{0}=10^{-38}$. This value is inversely proportional to the mass of the ultra-compact object. The behavior of the escape condition for BH and ABH without EH shows a similar trend of the required disk luminosity diverging as the position of the particle reaches the ISCO, as shown in the Figs.2(a) & 2(b). However, as the disk extends to the central ultra compact object in ABH without EH, the divergence of the normalized luminosity occurs in the ultra strong gravity region compared to the Schwarzschild black hole where this divergence occur at the ISCO, which is very far from the EH. It is evident from Eqns. (18) & (19), when the disk luminosity is low due to the low accretion rate, the wind generation from the BH is not possible. However, even with the very low accretion rate, the ABH without EH will be able to generate the winds due to their extremely small value of the ISCO and hence the large value of $\Gamma_{d}$. The Eqn.(18) shows that the $\Gamma_{d(ABH~{}without~{}EH)}\sim 3\times 10^{38}\Gamma_{b(BH)}$ for the same rate of accretion. Hence, the winds generated from the accretion disk around the ABH without EH will be much stronger compared to those generated from the disk around a BH. These calculations neglect general relativistic effects; however, they offer a rough estimate of the relative ease of wind generation from ABHs without an event horizon compared to BHs, and consequently, their role in the quenching of star formation.
@@ IV.5 Outflows from Naked Singularities
Further insight about the generation of outflows or jets from the naked singularities can be adduced via hydrodynamical simulations. Recently, researchers have performed hydrodynamical simulations of accretion onto Reissner-Nordström(RN) naked singularities and compared them with RN black holes[78]. The results were completely dependent on the existence of the event horizon, which governs the accretion pattern. That was ensured by the consideration of electrically neutral fluid, which will be independent of the amplitudes of charges of RN spacetimes under study. The typical behavior of the black hole event horizon is observed. All the matter being accreted will be absorbed by the event horizon; hence, the outflows from the black holes can only be generated with the presence of a magnetic field, radiation, or rotation. However, RN naked singularities were observed to be favorable for the formation of outflows and jets. It was observed from the simulations that the fluid can have mildly relativistic velocities having turned around the RN naked singularities. This initial simulations of naked singularities clearly show the distinctive behavior of existence and non-existence of an event horizon. The mere fact that the accretion disk extends to the very central region of astrophysical black holes makes them favorable for generating outflows or jets, depending on the efficiency discussed in the previous subsection.
@@ IV.6 Accelerating Winds
Recently, a significant increase in the outflow velocity with time was detected in the spectrum of the quasar SBS $1408+544$ ($z=2.337$)[79]. Using $\sim 130$ spectra acquired over 8 years, as part of the spectroscopic monitoring during the SDSS, these authors have determined that the CIV broad absorption line (BAL) is not only varying in strength but also systematically shifting towards higher velocities. The net velocity shift is $683$ $kms^{-1}$, corresponding to a mean acceleration of $1.04$ $cms^{-2}$ (however, the acceleration varied throughout the monitoring period). They consider a range of possible explanations for the observed velocity shift and argue that it is most likely explained either as a geometric effect, or as a real acceleration of the outflow, due to radiation pressure. In the second scenario, the observed acceleration of the BLR could be due to an increased speed of the outflow driven by a varying incident (ionizing) radiation from the disc[80]. Thus, they [79] estimate that a certain combination of BH mass and disk wind launch radius can reproduce the observed BLR acceleration. However, in this model, to have the acceleration increasing with time, either the wind originates predominantly at radii $r<1.65r_{L}$, since this model allows acceleration to increase only at such small radii, or the wind’s terminal velocity is changing with time[79]. These constraints arise mainly due to the comparatively low efficiency of BHs in generating radiation from in-falling matter. In contrast, an ABH can easily explain this observation because its radiative efficiency increases with time and asymptotically approaches $100\%$, as discussed in the previous section. The increasing radiative efficiency is expected to accelerate the winds, exerting an increasing amount of radiation pressure. A detailed modelling of this construct, supported by numerical simulation of ABHs, is needed.
@@ IV.7 Some Observational Signatures
Maximally rotating black holes can emit energy, fueled by their rotation, which can be several powers of ten above the Eddington rate for stellar-mass black holes, the typically observed power being around $10^{43.5}$ $erg/s$, apparently independent of BH mass [81]. This suggests that many young massive stars turn into maximally rotating BH[82, 83, 84] for a short time (about $10^{4}$ yr for a 10 solar mass BH, $M_{BH,1}$, with the time scales rising linearly with BH mass) period. For stellar mass BHs, as this energy output exceeds the Eddington luminosities by many powers of ten, their impact on the environment could be very substantial, assuming a sufficiently high star formation rate. This observed power matches Eddington luminosity only for the smallest SMBHs. Such estimates have been published for over 25 years (e.g., [86, 85]). In such a scenario, the star forming process itself does the required quenching, via freshly created BHs, and in case the initial rotation is higher than the limit allowed for freshly formed stellar-mass BHs, the disruption caused due to the feedback would be even more pronounced. This would indicate that freshly formed stellar mass BHs are not BHs with an event horizon, but they are ABHs (which would take about $10^{4}$ $M_{BH,1}$ years to be covered within an event horizon).
Referring to the observations, the magnetic fields observed in terms of $(B\times r)$ $=$ $10^{16.0\pm 0.12}$ $(G\times cm)$ in a wind that resembles the Parker wind[87] and carries powerful electric currents[88] could actually be a (short-lived) signal for the putative transition from a compact object without a horizon to a BH with a horizon.

### v3

@@ I Introduction
In this paragraph, we outline the basic picture set out in the present study. The long-term quenching of galaxies is primarily attributed to a dense, compact object situated at the galactic nucleus[41]. Conventional black holes possess event horizon, limiting matter accretion up to a boundary which is set at the “Innermost stable Circular Orbit” (ISCO) and the consequent upper bound on the generation of ultra-high-energy particles. The cosmic censorship conjecture[42, 43] remains a foundational yet unproven and not even properly formulated, assumption in black hole physics. While it holds in simplified models such as spherically symmetric dust collapse, a wide range of studies on more general collapse scenarios—including those involving inhomogeneous matter fields, pressure, and rotation—have shown that naked singularities can emerge as legitimate solutions to Einstein’s equations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Despite multiple attempts, no rigorous mathematical proof exists for any version of cosmic censorship, and the conjecture’s vague formulation continues to hinder progress. In this light, the theoretical exclusion of naked singularities is no longer viable. Since the formation of such singularities is a generic outcome in many physically reasonable collapse models, the focus now necessarily shifts toward their observational signatures. Recent studies have explored the possibility that some astronomical observations, including the first image of Sagittarius $A^{*}$, can be explained by naked singularity models, with the JMN-1 model [60] recognized as a leading black hole mimicker for Sgr $A^{*}$ [61]. Other works have also considered naked singularities as valid candidates [62, 63]. Distinguishing naked singularities from black holes observationally is crucial and can be approached using various methods, including analyzing shadows-which may resemble black holes or appear as full-moon images depending on metric parameters [64, 65], perihelion precession, where naked singularities show both positive and negative precession unlike black holes [66, 67, 68], and Lense-Thirring precession effects on orbiting gyroscopes or pulsar beams, which differ near naked singularities compared to black holes [69, 70]. Additional techniques include measurements of relativistic time delays of pulsar signals [71] and tidal forces [72], all of which offer promising routes to verify the existence of naked singularities.
In this work, we consider the astrophysical implications of naked singularities—viewed as ultra-compact, massive objects without event horizons—and explore their potential impact on galaxy evolution, particularly in comparison to traditional black holes. To address this, we propose here models of “astrophysical black holes” (ABHs) that are devoid of an event horizon (EH), hence characterized by a huge amount of matter confined within an extremely compact region, a situation which permits the disk of accreting matter to extend all the way to the centre and yield extremely high energy densities and temperatures. This access to the ultra-strong gravity regime facilitates generation of extremely energetic particles, thus contributing to the long-term quenching of star formation in the host galaxy and its halo region. Additionally, we introduce models of “stellar-mass astrophysical black holes” (StMABHs) which can arise from gravitational collapse of massive stars once they run out of their internal nuclear fuel. These StMABHs, prior to getting enveloped by the event horizon, create high-density regions capable of emitting very high-energy particles over short time scales. Such a situation may lead to short-term episodes of (localised) intra-galactic quenching. Potential observational signatures associated with the StMABHs are also pointed out here. Section II provides a brief overview of the recent observations related to this phenomenon, made with the James Webb Space Telescope (JWST). Section III touches upon the role SMBHs play in the galaxy quenching process. Section IV presents some salient features of our proposed models involving astrophysical black holes (ABH), in the context of long- and short-term quenching. Our conclusions are summarised in Section V.
@@ IV.1 A plausible scenario for long-term quenching
$ds^{2}=-e^{4\int_{0}^{\omega}\omega\frac{R^{\prime}}{R}dr}dt^{2}+\frac{R^{\prime 2}}{b(r)e^{4\int_{v}^{1}\frac{\omega}{v}dv}}dr^{2}+R^{2}d\Omega^{2}.$ (1)
@@
$ds^{2}=-e^{-2\int_{0}^{r}\frac{p^{\prime}}{\epsilon+p}dr}dt^{2}+\frac{R^{\prime 2}}{b(r)e^{2\int_{r}^{R}\frac{v^{\prime}}{R^{\prime}}dR}}dr^{2}+R^{2}d\Omega^{2}.$ (2)
@@ IV.3 The Role of Radiative and Mechanical Efficiency
Black holes with super-Eddington accretion would operate in a radiatively efficient, ‘quasar’ mode. In contrast, BHs with highly sub-Eddington accretion operate in a radiatively inefficient, ‘radio’ mode. Despite the lower radiative efficiency, this ‘radio’ mode generates substantial mechanical luminosity, driving winds and/or collimated jets[108]. The magnitude of the Eddington ratio determines the transition between the two accretion (i.e., feedback) modes. We recall the simplified treatment by Silk and Rees[86], where they estimated the effect of the accretion disk winds on the galaxy. The galaxy was assumed to be an isothermal sphere having a gas fraction, $f_{gas}$ and a density $\rho=v^{2}/2\pi Gr^{2}$, where $v$ is the velocity dispersion. As discussed in [86], the wind can propel the interstellar gas out of the galaxy with a constant velocity,
@@
Here, the wind luminosity is the fraction $f_{w}$ of the Eddington luminosity. The fraction $f_{w}$ is dependent on the radiation efficiency $\epsilon$ as[86],
@@
where $\gamma\equiv f_{gas}/32\pi^{3}f_{w}$. In theory then, a black hole could expel all the material out of its host galaxy, once its mass surpasses $\sim 10^{7}M_{\odot}$(for $\gamma\sim 1$ & $f_{w}\sim 0.01$)(however, these arguments hold for early galaxies).
@@
$\mathcal{F}(r)=-\frac{\dot{m}}{4\pi\sqrt{-g}}\frac{\omega^{\prime}}{(E-\omega L)^{2}}\int^{r}_{r_{in}}(E-\omega L)L^{\prime}dr.$ (12)
@@
A more practical measure is the luminosity, i.e. the energy that reaching an observer at infinity $\mathcal{L}_{\infty}$, per unit time. By integrating $d\mathcal{L}_{\infty}/dlnr=4\pi r\sqrt{-g}E\mathcal{F}$ with lnrlnr, we can calculate the net luminosity $\mathcal{L}_{\infty}$ observed at infinity. As the disc is truncated at the ISCO in the BH case, the efficiency is only $\sim 6\%$ as $\epsilon=1-E_{ISCO}=\mathcal{L}_{\infty}/\dot{m}c^{2}$. In the case of ABHs (see, JMN1 & JMN2), the accretion disk can extend up to the central high-density entity. However, the singularity is not yet formed, so the accretion disk extends up to the boundary of the central ultradense object, and the energy of the innermost circular orbit will then always be lesser than the $E_{ISCO}$. This essentially means that, as the collapse proceeds, the radiative efficiency of the ABH rises, which can be verified by calculating $\mathcal{L}_{\infty}$ for the dynamic metrics given in (1) & (2) using Eqns.(7)-(13). When the singularity has formed in the asymptotic time, the stable circular orbits can extend right up to the singularity. The radiative efficiency will then approach $100\%$. The $\gamma$ of the Eqn. (5) will be $\sim 17$ in that case, and we require $17$ times more mass than the typical BH, in order to expel all the material from the host galaxy. This estimate is meant for the feedback due to mechanically generated winds from the accretion disk. The calculation assumes that the mechanical efficiency is reversibly proportional to the radiative efficiency. The radiative efficiency of ABH being very high, it would have a lower efficiency for mechanical energy injection and hence required to be more massive. However, when the winds are accelerated by the disc radiation, the ABH will be able to generate high-energy winds with greater efficiency. The next subsection discusses this in detail with the example of the hydrodynamical simulation for a naked singularity spacetime. In the BH case, the radiation-dominated (quasar) mode sets in only when the accretion rate becomes very high. In contrast, an ABH can produce very efficient radiation-driven winds even when the accretion rate is very low.
@@ IV.4 Wind Generation
with ionization rate $\chi_{e}=1$, Eddington luminosity $L_{E}=4\pi GcMm_{p}/\sigma_{T}$ (here $m_{p}$ is the proton mass), and the disk luminosity[89],
@@
In the units of $2GM/c^{2}$, the ISCO for the Schwarzschild black hole is located at $r_{ISCO}=3$. However, the ISCO for the ABH without EH models discussed in the previous section would be very close to the central ultra-dense object. If we assume the ultra compact object to be of the size of the order of Planck length, then ISCO for a few solar mass ABH without EH will be at $r_{ISCO}=10^{-38}$. This value is inversely proportional to the mass of the ultra-compact object. The behavior of the escape condition for BH and ABH without EH shows a similar trend of the required disk luminosity diverging as the position of the particle reaches the ISCO, as shown in the Figs.2(a) & 2(b). However, as the disk extends to the central ultra compact object in ABH without EH, the divergence of the normalized luminosity occurs in the ultra strong gravity region compared to the Schwarzschild black hole where this divergence occur at the ISCO, which is very far from the EH. It is evident from Eqns. (18) & (19), when the disk luminosity is low due to the low accretion rate, the wind generation from the BH is not possible. However, even with the very low accretion rate, the ABH without EH will be able to generate the winds due to their extremely small value of the ISCO and hence the large value of $\Gamma_{d}$. The Eqn.(18) shows that the $\Gamma_{d(ABH~without~EH)}\sim 3\times 10^{38}\Gamma_{b(BH)}$ for the same rate of accretion. Hence, the winds generated from the accretion disk around the ABH without EH will be much stronger compared to those generated from the disk around a BH. These calculations neglect general relativistic effects; however, they provide a rough estimate of the relative ease of wind generation from ABHs without an event horizon compared to BHs, and consequently, their role in quenching star formation. Particle dynamics in the ultra-strong gravitational regime near ABH cannot be accurately modeled using the Newtonian framework applied in wind generation analyses. Although such approaches may provide approximate estimates, a precise treatment necessitates general relativistic numerical simulations, which we intend to pursue in future work. Previous hydrodynamical and general relativistic magnetohydrodynamic (GRMHD) studies of select naked singularity models have been conducted by other researchers [90, 91], and these works are reviewed in the following section.
@@ IV.5 Outflows from Naked Singularities
Further insight about the generation of outflows or jets from the naked singularities can be adduced via hydrodynamical simulations. Recently, researchers have performed hydrodynamical simulations of accretion onto Reissner-Nordström(RN) naked singularities and compared them with RN black holes[90]. The results were completely dependent on the existence of the event horizon, which governs the accretion pattern. That was ensured by the consideration of an electrically neutral fluid, which will be independent of the amplitudes of charges of RN spacetimes under study. The typical behavior of the black hole event horizon is observed. All the matter being accreted will be absorbed by the event horizon; hence, the outflows from the black holes can only be generated with the presence of a magnetic field, radiation, or rotation. However, RN naked singularities were observed to be favorable for the formation of outflows and jets. It was observed from the simulations that the fluid can have mildly relativistic velocities having turned around the RN naked singularities. This initial simulations of naked singularities clearly show the distinctive behavior of existence and non-existence of an event horizon. The mere fact that the accretion disk extends to the very central region of astrophysical black holes makes them favorable for generating outflows or jets, depending on the efficiency discussed in the previous subsection. However, the dynamics of real galaxies are governed by rotation, magnetic fields, and complex gas physics. Crucial astrophysical phenomena, such as jet launching, are fundamentally linked to both rotation and magnetic fields. Recent studies have employed general relativistic magnetohydrodynamic (GRMHD) simulations to investigate the accretion dynamics around JMN naked singularities [91]. These studies reveal that such naked singularities possess a centrifugal barrier that obstructs the direct accretion of rotating matter. Despite a reduction in magnetization within the funnel region, these singularities are capable of generating jet powers comparable to those observed in black holes. Moreover, strong outflows or jets emerge, with driving mechanisms that may vary with the specific spacetime geometry and spatial location—being dominated at times by gas pressure and at others by magnetic pressure. Importantly, the energy powering these outflows or jets is derived from the gravitational potential energy of the matter accumulated in the deep potential well. For particular parameter regimes, the JMN singularity can even produce jets exceeding the strength of those observed from black holes.
@@ IV.6 Accelerating Winds
Recently, a significant increase in the outflow velocity with time was detected in the spectrum of the quasar SBS $1408+544$ (z=2.337z=2.337)[92]. Using $\sim 130$ spectra acquired over 8 years, as part of the spectroscopic monitoring during the SDSS, these authors have determined that the CIV broad absorption line (BAL) is not only varying in strength but also systematically shifting towards higher velocities. The net velocity shift is $683$ $kms^{-1}$, corresponding to a mean acceleration of $1.04$ $cms^{-2}$ (however, the acceleration varied throughout the monitoring period). They consider a range of possible explanations for the observed velocity shift and argue that it is most likely explained either as a geometric effect, or as a real acceleration of the outflow, due to radiation pressure. In the second scenario, the observed acceleration of the BLR could be due to an increased speed of the outflow driven by a varying incident (ionizing) radiation from the disc[93]. Thus, they [92] estimate that a certain combination of BH mass and disk wind launch radius can reproduce the observed BLR acceleration. However, in this model, to have the acceleration increasing with time, either the wind originates predominantly at radii $r<1.65r_{L}$, since this model allows acceleration to increase only at such small radii, or the wind’s terminal velocity is changing with time[92]. These constraints arise mainly due to the comparatively low efficiency of BHs in generating radiation from in-falling matter. In contrast, an ABH can easily explain this observation because its radiative efficiency increases with time and asymptotically approaches $100\%$, as discussed in the previous section. The increasing radiative efficiency is expected to accelerate the winds, exerting an increasing amount of radiation pressure. A detailed modelling of this construct, supported by numerical simulation of ABHs, is needed.
@@ IV.7 Some Observational Signatures
Maximally rotating black holes can emit energy, fueled by their rotation, which can be several powers of ten above the Eddington rate for stellar-mass black holes, the typically observed power being around $10^{43.5}$ $erg/s$, apparently independent of BH mass [94]. This suggests that many young massive stars turn into maximally rotating BH[95, 96, 97] for a short time (about $10^{4}$ yr for a 10 solar mass BH, $M_{BH,1}$, with the time scales rising linearly with BH mass) period. For stellar mass BHs, as this energy output exceeds the Eddington luminosities by many powers of ten, their impact on the environment could be very substantial, assuming a sufficiently high star formation rate. This observed power matches Eddington luminosity only for the smallest SMBHs. Such estimates have been published for over 25 years (e.g., [99, 98]). In such a scenario, the star forming process itself does the required quenching, via freshly created BHs, and in case the initial rotation is higher than the limit allowed for freshly formed stellar-mass BHs, the disruption caused due to the feedback would be even more pronounced. This would indicate that freshly formed stellar mass BHs are not BHs with an event horizon, but they are ABHs (which would take about $10^{4}$ $M_{BH,1}$ years to be covered within an event horizon). Here we are referring to the ABH discussed in sec.IV.2. In [100], the final structural configuration of a massive star immediately prior to its collapse and the subsequent formation of a rotating black hole is presented. The portion of the stellar core destined to form the black hole possesses approximately $60\%$ more angular momentum than the theoretical upper limit. We propose that, during the collapse, a transient naked singularity phase may occur, allowing the system to shed the excess angular momentum. To further assess this possibility, one would ideally investigate gravitational collapse scenarios through detailed numerical simulations and examine the quantum gravity effects expected in the vicinity of naked singularities. Such studies could help clarify the nature of radiation or matter that may escape from these objects and determine whether a connection with the observed galactic winds can be established. Previous [101] work has explored quantum gravity effects near naked singularities, particularly within the framework of loop quantum gravity. These studies indicate that quantum corrections introduced at the final stages of gravitational collapse can generate a strong outward energy flux. However, these analyses have so far been limited to specific collapse models, such as those involving scalar fields, and have not yet been extended to more general astrophysically relevant matter configurations. At present, the absence of a complete and experimentally validated theory of quantum gravity prevents a fully rigorous investigation along these lines. Nevertheless, the scenario considered here may still be examined within an astrophysical context. In particular, one may search for possible quantum gravity signatures in the properties of these winds, which could eventually contribute to a phenomenological understanding of quantum gravity effects.
Referring to the observations, the magnetic fields observed in terms of $(B\times r)$ == $10^{16.0\pm 0.12}$ $(G\times cm)$ in a wind that resembles the Parker wind[102] and carries powerful electric currents[103] could actually be a (short-lived) signal for the putative transition from a compact object without a horizon to a BH with a horizon. However, this remains a preliminary proposal; further progress in this direction requires the explicit development of a detailed model followed by rigorous fitting to observational data.